passes through (5,-2) perpendicular to -8x + 9y =-22
Answer by mukhopadhyay(490)

Line perpendicular to -8x + 9y =-22 will have a slope of -9/8.
Equation of the line perpendicular to given line is y = (-9/8)x + k
Passes thru (5,-2):
y = (-9/8)x + k
=> -2 = (-9/8)5 + k
=> k = -2 + 45/8 = 29/8
Equation of the desirable line in slope intercept form is:
y = (-9/8)x + 29/8